# LLM-Powered Agents for Navigating Venice’s Historical Cadastre

### v1

@@ Abstract
Cadastral data reveal key information about the historical organization of cities but are often non-standardized due to diverse formats and human annotations, complicating large-scale analysis. We explore as a case study Venice’s urban history during the critical period from 1740 to 1808, capturing the transition following the fall of the ancient Republic and the Ancien Régime. This era’s complex cadastral data, marked by its volume and lack of uniform structure, presents unique challenges that our approach adeptly navigates, enabling us to generate spatial queries that bridge past and present urban landscapes. We present a text-to-programs framework that leverages Large Language Models (LLMs) to translate natural language queries into executable code for processing historical cadastral records. Our methodology implements two complementary techniques: a text-to-SQL approach for handling structured queries about specific cadastral information, and a text-to-Python approach for complex analytical operations requiring custom data manipulation. We propose a taxonomy that classifies historical research questions based on their complexity and analytical requirements, mapping them to the most appropriate technical approach. This framework is supported by an investigation into the execution consistency of the system, alongside a qualitative analysis of the answers it produces. By ensuring interpretability and minimizing hallucination through verifiable program outputs, we demonstrate the system’s effectiveness in reconstructing past population information, property features, and spatiotemporal comparisons in Venice.
@@
open-ended learning, developmental AI, language, culture, socio-cultural situatedness, reinforcement learning, autotelic agents, cognitive tools, natural language processing, large language models, culture models
@@ 1.1 Context
Historical cadastral records, widely distributed throughout Europe, serve as invaluable documents to reconstruct past urban and territorial information [1]. These records document property ownership, usage functions, and other essential elements for taxation, offering high confidence in their reliability due to their administrative purpose [2]. Often paired with cartographic mappings, these dual systems combine textual descriptions with geographic representations following standardized visual and ontological codes to minimize subjective interpretation and enhance utility for taxation [3]. The evolution of cadastral sources has been extensively studied, with analyses spanning specific case studies to comparative frameworks. Associated cartography, particularly before and after the Napoleonic introduction of geometric-parcel cadastres, reflects a shift towards standardized cartographic practices [4]. These sources are critical for reconstructing historical population data, property functions, and urban spatial dynamics.
These historical cadastral records offer rich opportunities for understanding urban development, social structures, and economic patterns. Researchers can investigate diverse questions ranging from property ownership dynamics (e.g., identifying dominant landholders or tracking wealth concentration) to spatial-functional analyses (e.g., mapping commercial activities in specific districts). While digitization has made these documents more accessible, the challenge lies in efficiently queryinh and analyzing this wealth of information to answer such research questions. Traditional manual methods of data extraction and analysis are time-consuming and limit the scale at which these historical insights can be pursued, creating a need for computational approaches that can systematically process these records while maintaining the nuanced understanding required for historical research.
To address these analytical challenges, this paper explores the application of Large Language Model (LLM)-based agents to analyze historical cadastre datasets, specifically focusing on two digitized Venetian cadastres: the 1740 Catastici from the Republic of Venice and the 1808 Sommarioni from the Napoleonic Kingdom of Italy [5] . These historical records present significant processing challenges, including orthographic variations, transcription errors, and non-standardized formats, as illustrated in Figure 1.
To effectively query the cadastral data, we propose a text-to-program approach that translates natural language queries into executable code. Within this strategy, we explore two complementary techniques. The first employs text-to-SQL translation, which is optimal for precise, structured queries about specific cadastral information (e.g., "Who owned property X in 1740?" or "What was the value of properties in district Y?"). The second utilizes text-to-Python translation, enabling more complex analytical operations and pattern recognition through custom code generation. This technique is particularly suitable for broader research questions requiring data manipulation and statistical analysis (e.g., "How did property ownership patterns change between 1740 and 1808?" or "What were the spatial distribution trends of different property types?").
@@
Figure 1: The role of LLMs in processing historical Cadastres. Processing historical records with orthographic variations and complex transcription details through text-to-SQL and text-to-Python approaches for systematic data analysis.
The remainder of this paper is structured as follows. In the next section, we present the cadastral data of Venice that forms the foundation of our study. Section 3 introduces a typology of historical questions relevant to this cadastral data, addressing different analytical needs and query complexities. Sections 5 and 6 detail our text-to-programs systems, demonstrating the implementation and capabilities of both text-to-SQL and text-to-Python approaches. We conclude by evaluating the accuracy and reliability of these systems, offering a typology of their error types and consistency metrics to ensure answer trustworthiness.
@@ 2 Cadastral data of the city of venice
The first geometric cadastre in Venice was established in 1808, adhering to French administrative standards [6]. As displayed in figure 2, it operates as a dual system, combining cartographic maps of parcels with textual records that document ownership, location, function, and area. Each parcel is assigned a unique number, which is cross-referenced with the records that catalog owners, toponyms, uses, and dimensions (see Figure 3b). Ownership records include individual names, family relationships, and institutions, while functions are classified using a codified Italian ontology from 1808 [7]. In particular, the terms reflect historical usage; for instance, a “shop” is referred to as ’bottega’, rather than the modern Italian term ’negozio’.
@@
Otherwise, the 1740 Catastico represents a textual survey system managed by the Collegio dei Dieci Savi at Rialto, designed to administer the Venetian tithe—a 10% property tax introduced in 1463. Property owners submitted self-declared ’Condizioni di Decima’ or ’Polizze’ which detailed property type, location, status, and income. These submissions were organized by district and sequentially numbered based on submission order, with taxation calculated from declared rents [8]. The overall informational structure of the document is displayed in figure 3a. Following a major archival fire in 1514, redecimation efforts occurred sporadically, with significant collections in 1514, 1537, 1566, 1581, 1661, 1711, and 1740. Unlike the 1808 cadastre, the Catastico did not integrate cartographic representation. Instead, records were generated through door-to-door surveys conducted by censors, who documented owners, tenants, toponyms, property functions, and rents (see Fig. 2-A).
These systems reflect complementary approaches to surveying. The 1808 cadastre is cartographic, parcel-based, and systematic, while the Catastico is textual, household-focused, and income-oriented. Despite these differences, both systems exhibit stable informational structures [8]. Both have been digitized and transcribed: in the 1808 cadastre, parcel identifiers, codes, and toponyms were automatically transcribed and subsequently verified manually [9, 10]. For the 1740 Catastico, geolocation was achieved by correlating toponyms with contemporary maps, reconstructing the censors’ survey paths, and identifying shared features between parcels recorded in 1740 and 1808.
@@
Together, these datasets encompass more than 34,000 data points, providing detailed information on owner professions, tenants, property functions, rents, areas, and geolocations.
@@ 3 A typology of historical questions related to cadastre data
In urban historical research, cadastral data are used to link people and urban functions to territories. There are several objectives when consulting these records. The first is to consult them to identify the location of one or more people or urban functions in a specific place or places. The second is to investigate more general principles and test hypotheses about specific places, groups of people, types of function, or compare periods of time, particularly the past with the present. These research questions often involve complex statistical operations to aggregate and compare information that spans multiple data points from potentially different datasets. In the first case we test the accuracy of LLM to find, combine and compute the information we need : the name(s) of owners, urban functions, toponyms and others. This allows LLM to be used as a historical dataset browser. In the second case, we try to understand the precision of LLM in answering complex questions that would detect a notion of semantic and temporal context of the historical dataset. We propose to organize the paper around those two distinct objectives in historical research browsing vs. prompting.
@@ 3.1 Browsing questions
In contrast, relational queries delve deeper into the dataset to examine the relationships and patterns among various data points, such as identifying how many families own properties across multiple categories. An example of such a question could be, “How many families own properties of more than one type category?“. These relational queries are essential for revealing trends in property ownership and usage, providing insights into socio-economic dynamics within urban settings. Table 3.1 displays an aggregation as well as a relational question with their related SQL queries. In section 5, we will evaluate the performance of a text2SQL agent on a set of 100 hand-crafted browsing questions about the Catastici 1740 dataset (provided in Appendix Section A.1).
@@
What is the total rent revenue generated from properties of the "bottega da casarol" variety? SELECT SUM("Rent_Income") FROM catastici WHERE "Property_Type" = "bottegadacasarol";
How many families own properties of more than one type category? SELECT COUNT(*) FROM ( SELECT "Owner_Family_Name" FROM catastici GROUP BY "Owner_Family_Name" HAVING COUNT(DISTINCT "Property_Type") > 1 ) AS families_with_multiple_types;
@@ 3.2 Prompting questions
After careful analysis of the datasets and their potential applications, we identified that meaningful questions about historical cadastral data could be organized into four distinct categories. The first category leverages the geocoordinates of the cadastre entries to examine spatial distributions, enabling queries that bridge past and present urban landscapes. By relating historical properties to relatively stable urban landmarks such as churches and squares (extracted from OpenStreetMap [11]), we can investigate how individuals and properties were allocated across diverse areas. Although these landmarks may have undergone modifications or reconstructions over time, they often maintain their general location and social function, serving as semipersistent spatial anchors for historical analysis. The second category is dedicated to building functions, exploring the intended purposes or uses of various structures within the urban environment. The third category focuses on personal information, examining demographic and socioeconomic characteristics associated with individuals in the cadastral data. Finally, the fourth category targets temporal analysis, specifically comparing data over two distinct periods to reveal trends, shifts, or patterns over time.
Table 2: Examples of prompting questions. alongside with their category and expected output format.
@@
Unlike browsing questions, designing SQL queries for prompting questions presents significant challenges due to their inherently complex nature. Prompting questions often require insights that extend beyond the information readily available in the dataset’s structure or columns. The intricate operations necessary for these questions move beyond simple data filtering and aggregation, involving advanced processes such as semantic searches, spatial computations, and statistical tests or correlation evaluations. Additionally, certain prompting questions demand the incorporation of external knowledge or common-sense reasoning, which cannot be encapsulated purely through SQL. In section 6, we introduce the text-to-Python agent, which transforms prompting questions into Python programs to effectively generate answers.
4 Approaches to Historical Cadastral Data Processing
Processing historical cadastral data is challenging due to irregular formats, orthographic variations, and complex historical annotations. Existing computational methods fall into three main categories: machine learning, rule-based approaches, and large language model (LLM)-based code generation.
Machine Learning Approaches: Supervised models like TabularNet [12] and TableFormer [13] have improved table parsing by leveraging neural architectures, while STab [14] introduced self-supervised learning for diverse tabular data. More recent methods, such as mixed-type tabular data synthesis [15], aim to reconstruct incomplete records. However, these approaches require extensive training data and struggle with the variability of historical sources.
Rule-Based Methods: These approaches rely on predefined extraction rules for normalizing toponyms, recognizing property descriptions, and applying historical ontologies. While interpretable, they lack adaptability to inconsistent datasets.
LLM-Based Code Generation: Instead of relying on pre-trained models for structured parsing, we leverage LLMs to dynamically generate executable queries—either SQL for structured retrieval or Python for complex analyses. This approach offers: (1) Flexibility – No need for labeled training data, making it adaptable to diverse historical datasets; (2) interpretability – Produces verifiable code rather than opaque model outputs; and (3) Scalability – Handles both simple lookups and complex spatial-temporal analyses.
In the following sections, we present two complementary approaches: a SQL-based system for precise data retrieval and a Python-based system for complex analytical tasks. These approaches demonstrate how code generation through LLMs can offer a more adaptable and maintainable solution for historical data analysis compared to traditional ML methods.
5 Browsing cadastre data with SQL agents
5.1 Method
Questions and data specifications. We develop a set of hundred questions designed to facilitate the exploration of the Catastici 1740 dataset. These questions are available in Section A.1. They were handcrafted to cover both retrieval and relational concepts (as described in Section 3.1) about the information contained in a simplified version of the Catastici Dataset including seven columns: Catastici_ID[integer], Owner_ID[integer], Owner_First_Name[text], Owner_Family_Name[text], Property_Type[text], Rent_Income[integer], Property_Location[text]
Model. To convert these natural language questions into SQL queries, we employ the open-source text-to-SQL model, CodeS-7B [16]. In accordance with the model’s specifications, each prompt includes both a comprehensive description of the table’s metadata and detailed information about its columns. We assess the model’s performance in both zero-shot and three-shot prompting scenarios, with the results detailed in the subsequent subsection. Each question is fed into the system a total of four times, and the SQL query selected is determined through a majority voting mechanism. The resulting SQL query is executed using SQLite to obtain the corresponding answers. For further clarity, examples of the prompts designed to interact with the CodeS-7B model are provided in Appendix A.2.

### v2

# LLM Agents for Interactive Exploration of Historical Cadastre Data: Framework and Application to Venice

@@ Abstract
Cadastral data reveal key information about the historical organization of cities but are often non-standardized due to diverse formats and human annotations, complicating large-scale analysis. We explore as a case study Venice’s urban history during the critical period from 1740 to 1808, capturing the transition following the fall of the ancient Republic and the Ancien Régime. This era’s complex cadastral data, marked by its volume and lack of uniform structure, presents unique challenges that our approach adeptly navigates, enabling us to generate spatial queries that bridge past and present urban landscapes. We present a text-to-programs framework that leverages Large Language Models (LLMs) to process natural language queries as executable code for analyzing historical cadastral records. Our methodology implements two complementary techniques: a SQL agent for handling structured queries about specific cadastral information, and a coding agent for complex analytical operations requiring custom data manipulation. We propose a taxonomy that classifies historical research questions based on their complexity and analytical requirements, mapping them to the most appropriate technical approach. This framework is supported by an investigation into the execution consistency of the system, alongside a qualitative analysis of the answers it produces. By ensuring interpretability and minimizing hallucination through verifiable program outputs, we demonstrate the system’s effectiveness in reconstructing past population information, property features, and spatiotemporal comparisons in Venice.
@@
Digital Humanities, Computational Urban Research, Large Language Models
[…]
Plain Language Summary
This study applies Large Language Models (LLMs) to historical cadastral data, focusing on Venice from 1740 to 1808, a period of major political and social change. Cadastral records, while valuable for urban history, are often unstandardized and challenging to analyze at scale. To address this, a framework is introduced that converts natural language queries into executable code, enabling structured database searches (SQL) and complex analytical operations (Python). By categorizing historical research questions based on complexity, the methodology ensures accuracy and minimizes errors. This approach reconstructs past urban landscapes, uncovering insights into property ownership, land use, and socio-economic patterns. The findings highlight LLMs’ potential for historical urban studies, offering a scalable and reliable method applicable across disciplines.
@@ 1.1 Context
Historical cadastral records, widely distributed throughout Europe, serve as invaluable documents to reconstruct past urban and territorial information kain_cadastral_1992 (18). These records document property ownership, usage functions, and other essential elements for taxation, offering high confidence in their reliability due to their administrative purpose bloch_plans_1929 (3). Often paired with cartographic mappings, these dual systems combine textual descriptions with geographic representations following standardized visual and ontological codes to minimize subjective interpretation and enhance utility for taxation bourguet_dechiffrer_1988 (4). The evolution of cadastral sources has been extensively studied, with analyses spanning specific case studies to comparative frameworks. Associated cartography, particularly before and after the Napoleonic introduction of geometric-parcel cadastres, reflects a shift towards standardized cartographic practices clergeot_cent_2007 (8). These sources are critical for reconstructing historical population data, property functions, and spatial dynamics, and offer rich opportunities for analyzing urban development and socio-economic structures.
While digitization has increased access to cadastral records, efficiently querying and analyzing them remains a significant challenge. Traditional methods—based on close reading and manual data extraction—are labor-intensive and difficult to scale across time and space. As researchers seek to engage with these sources at scale, there is a growing need for computational approaches that preserve historical nuance while enabling forms of distant reading. This also aligns with recent calls for more design-oriented methodologies in cadastral research, which emphasize the development of flexible tools for information processing CAGDAS201177 (5).
1.2 Objectives and Contributions
This study introduces a generalizable framework that uses Large Language Models (LLMs) to assist historians in querying and analyzing historical cadastral data. Our objectives are threefold:
  • To overcome the structural heterogeneity of historical cadastral sources, which include orthographic variations, transcription inconsistencies, and non-standardized formats that hinder large-scale analysis (see Figure 1);
  • To enable domain experts to interact with these datasets using natural language, by processing research questions as executable programs that retrieve, aggregate, or analyze the relevant data;
  • To evaluate the reliability, consistency, and scope of LLM-based agents when applied to diachronic and spatially grounded historical records.
Our main contributions are:
  • A typology of historical research questions tailored to cadastral analysis, which distinguishes between browsing (structured lookups) and prompting (contextual or multi-dataset queries), and maps them to the appropriate processing method.
  • A dual-agent system for historical data analysis, combining two complementary strategies:
    • A SQL agent for structured, idiographic queries;
    • A Coding agent for more complex analytical operations, such as spatiotemporal comparisons and pattern detection.
  • A scalable and interpretable workflow that outputs verifiable code. This ensures that each answer is grounded in the source data and allows researchers to inspect or modify the generated analysis logic.
  • A case study on the cadastres of Venice (1740 and 1808), through which we evaluate our framework on real-world archival data. This includes an open benchmark of 240 expert-curated research questions (100 for structured browsing and 140 for complex prompting) that span spatial, functional, personal, and temporal analyses. The benchmark is designed to reflect authentic urban historical research and support future comparative evaluations. Two example queries are shown in Figure 1.
@@ 1.2 Objectives and Contributions
Figure 1: The role of LLMs in processing historical Cadastres. Processing historical records with orthographic variations and complex transcription details through SQL and coding agents for systematic data analysis.
1.3 Related Work
Processing historical cadastral data requires bridging challenges related to unstructured formats, orthographic inconsistency, and diachronic language variation. Existing computational strategies fall broadly into three categories: supervised machine learning models, rule-based methods, and more recently, LLM-based code generation.
Machine learning approaches have leveraged advances in tabular modeling to support structured data extraction and reasoning. Models such as TabularNet du2021tabular (12) and TableFormer yang-etal-2022-tableformer (32) integrate neural architectures for improved table parsing, while STab hajiramezanali2022stab (16) introduces self-supervised learning for diverse tabular data. Other methods, such as those presented in zhang2024mixedtype (33), aim to reconstruct incomplete historical records through data synthesis. However, these approaches generally require large labeled datasets and may struggle with the variability of historical formats.
Rule-based methods are often designed with an assumed underlying relational structure SHIGAROV2017123 (29), and rely on predefined rules for normalizing toponyms garbin2005disambiguating (14), recognizing descriptions, and applying historical ontologies. While interpretable and often robust in controlled settings, they tend to be brittle when faced with noisy, ambiguous, or non-standard historical data.
LLM-based code generation represents a more flexible alternative. Instead of relying on fixed schema or extensive training, these systems generate executable code—typically SQL for structured lookups or Python for custom analyses—directly from user queries. This paradigm has multiple advantages: (1) it eliminates the need for extensive labeled training data, (2) it produces verifiable outputs through executable programs, and (3) it scales across a wide range of query complexities, from simple lookups to advanced spatial-temporal operations. Building on recent advances in LLM-based data agents, our framework adapts and extends developments in (1) robustness, (2) flexibility, (3) modularity, and (4) adaptability to the specific challenges of historical cadastral analysis.
(1) Robustness is addressed through InfiAgent-DABench hu_infiagent-dabench_2024 (17), which formalizes benchmarks for execution consistency—defined as the statistical stability of results from repeated executions of equivalent queries. We adopt this methodology to quantify run-to-run variance as a proxy for output reliability.
(2) Flexibility is tackled by OpenAgent xie_openagents_2023 (31), which integrates SQL and Python execution within a single agentic framework, enabling seamless transitions between structured retrieval and procedural computation. While this architecture informed our design, our experiments confirm that SQL alone is ill-suited to the irregular schemas and semantic variability characteristic of cadastral sources. Kapoor et al. kapoor_ai_2024 (19) further propose techniques for improving cost efficiency and reducing debugging iterations, which have guided our implementation choices.
(3) Modularity is exemplified by CodeChain le_codechain_2024 (20), which promotes reusable, composable analysis components—an essential property when applying similar transformations across heterogeneous datasets.
(4) Adaptability is central to KwaiAgents pan_kwaiagents_2024 (27), whose architecture combines planning, memory, and tool use to support dynamic task execution in varied contexts. We draw on this principle to develop mechanisms that select and combine spatial, temporal, and semantic analyses according to query context and dataset properties as in Majumder et al. majumder_data-driven_2024 (24).
@@ 2 Cadastral Data of the City of Venice
The 1740 Catastico represents a textual survey system managed by the Collegio dei Dieci Savi in Rialto, designed to administer the Venetian tithe, a 10% property tax introduced in 1463. Property owners submitted self-declared ’Condizioni di Decima’ or ’Polizze’ which detailed property type, location, status, and income. These submissions were organized by district and sequentially numbered based on submission order, with taxation calculated from declared rents chauvard_catastici_2007 (7). The general information structure of the document is displayed in Figure 3a. Following a major archival fire in 1514, redecimation efforts occurred sporadically, with significant collections in 1514, 1537, 1566, 1581, 1661, 1711, and 1740. Unlike later geometric cadastres (e.g., the first one introduced in Venice in 1808), the Catastico did not integrate cartographic representation. Instead, records were generated through door-to-door surveys conducted by censors, who documented owners, tenants, toponyms, property functions, and rents (see Fig. 2-A).
@@
In contrast, the first geometric cadastre in Venice was established in 1808 pavanello_i_1981 (28), adhering to French administrative standards clergeot_recueil_2008 (9). As displayed in Figure 2, it operates as a dual system, combining cartographic maps of parcels with textual records that document ownership, location, function, and area. Each parcel is assigned a unique number, which is cross-referenced with the records that catalog owners, toponyms, uses, and dimensions (see Figure 3b). Ownership records include individual names, family relationships, and institutions, while functions are classified using a codified Italian ontology from 1808 di_lenardo_approche_2021 (10). In particular, the terms reflect historical usage; for instance, a “shop” is referred to as ’bottega’, rather than the modern Italian term ’negozio’.
These systems reflect complementary approaches to surveying. The 1808 cadastre is cartographic, parcel-based, and systematic, while the Catastico is textual, household-focused, and income-oriented. Despite these differences, both systems exhibit stable informational structures chauvard_catastici_2007 (7). Both have been digitized and transcribed: in the 1808 cadastre, parcel identifiers, codes, and toponyms were automatically transcribed and subsequently verified manually ares_oliveira_machine_2017 (1, 2). For the 1740 Catastico, geolocation was achieved by correlating toponyms with contemporary maps, reconstructing the censors’ survey paths, and identifying shared features between parcels recorded in 1740 and 1808.
@@
Together, these datasets encompass more than 34,000 data points, providing detailed information on owner professions, tenants, property functions, rents, areas, and geolocations. In this context, a ’data point’ corresponds to digitization of a single cadastral entry (row) containing all associated attributes such as parcel identifier, owner, profession, tenant and function with occasional missing values across fields.
@@ 3 A typology of historical questions related to cadastre data
In urban historical research, cadastral data are used to link people and urban functions to territories Ekamper15032010 (13, 2, 21). There are several objectives when consulting these records. The first is to consult them to identify the location of one or more people or urban functions in a specific place or places. The second is to investigate more general principles and test hypotheses about specific places, groups of people, types of function, or compare periods of time, particularly the past with the present. These research questions often involve complex statistical operations to aggregate and compare information that spans multiple data points from potentially different datasets.
In our evaluation, the first type of question assesses an LLM ’s ability to retrieve, combine, and compute relevant details—such as owners’ names, urban functions, and toponyms—thus using the model as a historical dataset browser. The second type probes the model’s capacity to answer more complex queries that require semantic understanding and spatio-temporal reasoning within the historical context.
@@ 3.1 Browsing questions
In contrast, relational queries delve deeper into the dataset to examine the relationships and patterns among various data points, such as identifying how many families own properties across multiple categories. An example of such a question could be, “How many families own properties of more than one type category?“. These relational queries are essential for revealing trends in property ownership and usage, providing insights into socio-economic dynamics within urban settings. Table 3.1 displays an aggregation as well as a relational question with their related SQL queries. The set of 100 hand-crafted browsing questions about the Catastici 1740 dataset are available in the Additional Methods.
@@
What is the total rent revenue generated from properties of the ‘bottega da casarol’ variety? SELECT SUM("Rent_Income") FROM catastici WHERE "Property_Type" = "bottegadacasarol"; How many families own properties of more than one type category? SELECT COUNT(*) FROM ( SELECT "Owner_Family_Name" FROM catastici GROUP BY "Owner_Family_Name" HAVING COUNT(DISTINCT "Property_Type") > 1 ) AS families_with_multiple_types;
@@ 3.2 Prompting questions
After careful analysis of the datasets and their potential applications, we identified that meaningful questions about historical cadastral data could be organized into four distinct categories. The first category leverages the geocoordinates of the cadastre entries to examine spatial distributions, enabling queries that bridge past and present urban landscapes. By relating historical properties to relatively stable urban landmarks such as churches and squares (extracted from OpenStreetMap OpenStreetMap (26)), we can investigate how individuals and properties were allocated across diverse areas. Although these landmarks may have undergone modifications or reconstructions over time, they often maintain their general location and social function, serving as semipersistent spatial anchors for historical analysis. The second category is dedicated to building functions, exploring the intended purposes or uses of various structures within the urban environment. The third category focuses on personal information, examining demographic and socioeconomic characteristics associated with individuals in the cadastral data. Finally, the fourth category targets temporal analysis, specifically comparing data over two distinct periods to reveal trends, shifts, or patterns over time.
Table 2: Examples of prompting questions. Alongside their category and expected output format.
@@
Unlike browsing questions, designing SQL queries for prompting questions presents significant challenges due to their inherently complex nature. Prompting questions often require insights that extend beyond the information readily available in the dataset’s structure or columns. The intricate operations necessary for these questions move beyond simple data filtering and aggregation, involving advanced processes such as semantic searches, spatial computations, and statistical tests or correlation evaluations. Additionally, certain prompting questions demand the incorporation of external knowledge or common-sense reasoning, which cannot be encapsulated purely through SQL.
4 Methods
4.1 Overview
Our framework combines two complementary approaches, each tailored to a category of research questions defined in Section 3:
  1. 1.
    SQL Agent — designed for browsing questions that involve structured lookups and relational queries over a single historical dataset.
    • Input: a natural language question.
    • Output: an SQL query executed on the target table to return results.
  2. 2.
    Coding Agent — designed for prompting questions requiring integration of multiple datasets and complex operations, such as spatial, temporal, or semantic reasoning.
    • Input: a natural language question.
    • Output: an executable Python program that performs the required analysis, optionally including multi-step planning, entity extraction, and consistency evaluation.
While both agents generate executable code directly from natural language, their architectures differ: the SQL Agent uses a single text-to-SQL model over a defined schema, whereas the coding agent coordinates multiple specialized LLM components to extract entities, plan analyses, and generate Python code.
4.2 Technical Data Representation
Our evaluation uses two historical cadastral datasets and a modern geographic reference dataset, represented in a simplified, analysis-ready form for the purposes of this study.
  • Catastici (1740) — a tabular dataset with seven columns: Catastici_ID [integer], Owner_ID [integer], Owner_First_Name [text], Owner_Family_Name [text], Property_Type [text], Rent_Income [integer], Property_Location [text].
  • Sommarioni (1808) — a comparable cadastral dataset with owner and property attributes, structured for cross-temporal comparison with the Catastici.
  • Landmarks dataset — a set of semi-persistent urban features (e.g., churches, squares) extracted from OpenStreetMap, represented as point geometries. These landmarks serve as spatial anchors for contextualizing historical property locations within a stable geographic framework.
4.3 SQL Agent
Questions. We curated 100 expert-designed browsing questions on the Catastici (1740) dataset, covering simple retrieval and relational queries, as detailed in the previous section